OpenStudy (anonymous):

Jacob has a piece of toast that has butter on one side, and he dropped it twice. ....

OpenStudy (anonymous):

Jacob has a piece of toast that has butter on one side, and he dropped it twice. Both times it landed with the butter side up. If he drops it two more times, what is the probability that it will have landed butter side up a total of three times? one over four one over two one over three two over three

OpenStudy (anonymous):

I would say 1/2

OpenStudy (anonymous):

How do you know? Can you say the steps plz?:)

geerky42 (geerky42):

Think of toast as coin. it already landed side up twice, so two times so far. And then Jacob decided to drop toast again. Let say side up is H and side down is T If we want it to land side up three times in total, we must have it land H one time and other times T. we have two possible outcomes; HT or TH out of four outcomes in total \(\rightarrow \dfrac{2}{4} = \boxed{\dfrac{1}{2}}\)

geerky42 (geerky42):

Is that clear?

OpenStudy (anonymous):

Yes, thx! Do you know how would I find the probability of this: Jim picked a card from a standard deck, looked at it, and then put it back. He then picked a second card. What is the probability that Jim picked a heart or an ace on either pick? one over fifty two two over fifty two sixteen over fifty two seventeen over fifty two *I don't see how 52 is in denominator..?

geerky42 (geerky42):

yeah, 52 is in denominator, and numerator is total number of heart and aces. there is 4 aces, right? and there is 13 hearts. Since one of ace is already heart, we don't want to count it twice, so numerator is either 4+12 or 3+13

geerky42 (geerky42):

52 is in denominator because there is 52 cards in total.

geerky42 (geerky42):

\( Probability = \dfrac{Number \space of\space ways\space it\space can \space happen }{Total \space number\space of \space outcomes}\)

OpenStudy (anonymous):

Oh, I see, so it would be 16/52! Thx again. I know nothing of how many cards are in a deck. lol

geerky42 (geerky42):

oh i see, that fine lol. just so you know, card of desk is make up of four suits (Club, Spade, Heart and Diamond) and each suit has 13 cards; ace, 2 to 10, then J, Q, and K.

geerky42 (geerky42):

13*4 = 52